The base of a mountain is at sea level where the gravitational field strength is 9.810N/kg. The value of the gravitational field strength at the top of the mountain is 9.790N/kg. Calculate the height of the mountain above the sea level.

1
Expert's answer
2021-02-22T10:27:30-0500

Answer

gravitational field strength at height h

"g_h=g_s(1-\frac{h}{2R})"

So height can be written as

"h=2R(1-\frac{g_h}{g_s})"

"h=2\times6400000\times(1-\frac{9.79}{9.810})"

h=26.1Km